OpenStudy (anonymous):

Use slopes to show that the points are vertices of the parallelogram. A(-3,1), B(5,3),C(3,0), D(-5,-2)

OpenStudy (anonymous):

find the slope of A-B, B-C, C-D, and D-A compare them, you should have 2 sets of parallel slopes perpendicular to one another

OpenStudy (anonymous):

I don't understand the part where you say they are perpendicular to one another...

OpenStudy (ajprincess):

Show that AB and CD are parallel and BC and AD are parallel.

OpenStudy (anonymous):

m(a,b)=1/4, m(b,c)=3/2, m(c,d)=1/4 m(d,a)=3/2

hero (hero):

Okay, so now use the definition of a parallelogram to finish your proof.

OpenStudy (anonymous):

Since m(a,b)=m(c,d) and m(b,c)=m(d,a) then the points are vertices of a parallelogram.

OpenStudy (anonymous):

How would I be able to solve a problem like this where my polygon is a right triangle, trapezoid or rectangle?

OpenStudy (ajprincess):

you need to knw the properties of each of them.using those properties u can solve.

OpenStudy (anonymous):

I saw it with the parallelogram and rectangle but not for a trapezoid and right triangle.